I need to solve this problem and I can'tfigure it out:
square root of x^2-15=square root of 2x

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

sqrt(x^2-15)=sqrt2x squaring both sides of the equation we get.
x^2-15=2x
x^2-2x-15=0
(x-5)(x+3)=0
x-5=0
x=5 answer.
x+3=0
x=-3 answer.

Answer by edjones(8007)

sqrt(x^2-15)=sqrt(2x)
square both sides: x^2-15=2x
x^2-2x-15=0
(x-5)(x+3)=0
x=-3 or x=5
Ed